Say you want to make a sling by swinging a mass M of 1.9 kg in a horizontal circle of radius 0.030 m, using a string of length 0.030 m. You wish the mass to have a kinetic energy of 14.0 Joules when released. How strong will the string need to be? (How much tension will there be in the string?)
Answer #1

Here Centripetal force = mv^2/r

= 0.5mv^2 *2/r

= KE*2/r

= 14*2/0.030 = 933.33 N

Gravitational force = mg = 1.9*9.8= 18.62 N

Tension required = sqrt(933.33^2+18.62^2)

= 933.5 N answer